OpenStudy (anonymous):

Use the quadratic formula to solve x2 - 2x - 3 = 0. x = -3 or x = 1 x = -1 or x = 3 x = 1 or x = -2 x = -2 or x = -1

OpenStudy (jdoe0001):

\(\bf \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -2}}x{\color{green}{ -3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im not understanding

OpenStudy (jdoe0001):

\(\bf \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -2}}x{\color{green}{ -3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ (-2)}} \pm \sqrt { {\color{red}{ (-2)}}^2 -4{\color{blue}{ (1)}}{\color{green}{ (-3)}}}}{2{\color{blue}{ (1)}}}\)

OpenStudy (anonymous):

x=1 and x=-2?

OpenStudy (jdoe0001):

well \({\bf \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -2}}x{\color{green}{ -3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ (-2)}} \pm \sqrt { {\color{red}{ (-2)}}^2 -4{\color{blue}{ (1)}}{\color{green}{ (-3)}}}}{2{\color{blue}{ (1)}}} \\ \quad \\ x=\cfrac{2\pm\sqrt{4+12}}{2}\implies x=\cfrac{2\pm\sqrt{16}}{2}\implies x=\cfrac{2\pm\sqrt{4^2}}{2} \\ \quad \\ x=\cfrac{\cancel{ 2 }\pm \cancel{ 4 }}{\cancel{ 2 }}\implies x= 1\pm 2}\to \large \begin{cases} x=1+2\to &3\\ x=1-2\to &-1 \end{cases}\)